The radius of the outer sphere of a spherical capacitor is five times the radius of its inner shell. What are the dimensions of this capacitor of its capacitance is 5.00 pF?


1
Expert's answer
2021-10-27T10:17:11-0400

The capacitance of a spherical capacitor is


"C=\frac{4\pi\epsilon_0}{1/a-1/b}=\frac{4\pi\epsilon_0}{1/a-1/5a},\\\space\\
a=\frac C{5\pi\epsilon_0}=0.036\text{ m},\\
b=5a=0.18\text{ m}."